OpenStudy (vortish):

factor the trinomial 6y^2-9y+5

OpenStudy (anonymous):

U sure that question u typed is correct

OpenStudy (anonymous):

I guess it means simplify

OpenStudy (vortish):

no its 6y^2 -9y +5

OpenStudy (anonymous):

first you need to multiply 6*5, which is 35. Now you need to find the factors of 35 that when added together will give you -9

OpenStudy (vortish):

6*5=30 not 35

OpenStudy (anonymous):

yes sorry

OpenStudy (vortish):

the only factors that i come up with are 3/10 = 30 and 6/5 = 30

OpenStudy (anonymous):

there are no factors that equal 30 and add to -9

OpenStudy (vortish):

not that i can figure out

OpenStudy (anonymous):

there isn't the only factors are -1*-30, -2*-15,-3*-10 and -5*-6

OpenStudy (vortish):

but even those do not come out to -9

OpenStudy (anonymous):

are you doing y-intercept

OpenStudy (vortish):

no just factoring that trinomial

OpenStudy (vortish):

Nburgher it is a prime trinomial ty you for your helop

OpenStudy (vortish):

gave you a medal to ty

OpenStudy (anonymous):

thank you your welcome